1. Mickey Mouse Company had zero units in its Work in Process Inventory. During the period it completed and transferred out 25,000 units and there were 5,000 units left in work in process inventory. The ending work in process were 100 % complete in materials and 40% complete as to conversion cost.

Required: i. How many units were in production during the period?

ii. Compute equivalent units for Materials and Conversion Cost under the Weighted Average Method.

iii. If total materials cost is $90,000 and total conversion cost is $71,000 at the end of the period.

Compute: a. The cost of goods completed and transferred out.

b. The total cost of ending work in process inventory.
Answer #1

i) Units in rpduction during period = 25,000 + 5,000 Units in rpduction during period = 30,000 ii) Materials Conversion costs

A B D E с Equivalent Units Physical units Materials Conversion costs 0 1 2 Quantities 3 Units to be accounted for 4 Work in p

A E B с D Equivalent Units Physical units Materials Conversion costs 0 1=25000+5000 =+B4+B5 25000 5000 =+B9*100% =+B9*100% =+